A project is expected to generate annual revenues of $119,300, with variable costs of $75,400, and fixed costs of $15,900. The annual depreciation is $3,950 and the tax rate is 34 percent. What is the annual operating cash flow?

Hint: Revenue - FC - VC - Depr. = EBIT. Taxes = EBIT x tax rate. OCF = EBIT + Depreciation - Taxes (same as chapter 2).

a. $61,143

b. $28,000

c. $19,823

d. $31,950

e. $45,243

THANKS IN ADVANCE!
Answer #1
Revenue 119300
Less:Variable cost 75400
Less:Fixed cost 15900
Less:Depreciation 3950
EBIT 24050

Taxes=EBIT*Tax rate

=24050*34%=$8177

OCF=EBIT+Depreciation-Taxes

=24050+3950-8177

=$19823.